Calculate the mass of oxalic acid dihydrate H2C2O4x2H2O
1
Expert's answer
2016-10-19T13:58:11-0400
Mr(H2C2O4x2H2O) = 2Ar(H) + 2Ar(C)+ 4Ar(O) + 2*2Ar(H) + 2Ar(O) = 2*1 + 2*12 + 4*16 + 4*1 + 2*16 = 126